# A data-driven method to estimate contamination from light ion beam transmutation at colliders

### v1

@@ Abstract
Collisions of relativistic light ions such as oxygen, neon, and magnesium, have been proposed as way to examine the system-size dependence of dynamics typically associated with the quark-gluon plasma produced in collisions of heavier ions such as xenon, gold, or lead. Recent efforts at both the Relativistic Heavy Ion Collider and Large Hadron Collider have produced large datasets of proton-oxygen, oxygen-oxygen, and neon-neon collisions, catalyzing intense interest in experimental backgrounds associated with light ion collisions. In particular, electromagnetic dissociation of light ions while they are circulating in a collider can result in beam contamination that is difficult to simulate precisely. Here we propose a data-driven method for evaluating the potential impact of beam contaminants on physics analyses. The method exploits the time-dependence and smaller size of contaminant ion species to define control regions that can be used to quantify potential contamination effects. A simple model is used to illustrate the method and to study its robustness. This method can inform studies of recent LHC and RHIC data and could also be useful for future light ion programs at the LHC and beyond.
@@ I Introduction
In the field of high energy nuclear physics, collisions of relativistic heavy ions, e.g., xenon, gold or lead, are used to study the emergent properties of the strong force. In these collisions, a strongly interacting hot medium, known as a quark-gluon plasma (QGP), is created Busza et al. (2018); Niida and Miake (2021); Harris and Müller (2024). The QGP exhibits many interesting and nontrivial dynamics including collective fluid-like behavior Romatschke and Romatschke (2019); Heinz and Snellings (2013) and results in strong modifications to particle production as compared to proton-proton collision baselines Khachatryan et al. (2017); Aad et al. (2015); Abelev et al. (2014).
@@
Perhaps surprisingly, critical examination of other collision systems having a much smaller size, such as proton-gold, deuteron-gold, helium-3-gold Aidala et al. (2018), and proton-lead collisions Aaboud et al. (2017); Khachatryan et al. (2015) have revealed similar behaviors as heavy ion collisions, prompting the question as to whether a QGP droplet is formed in these collisions. To try to gain more insight into this question, collisions of light ions such as oxygen, neon, and magnesium, have been proposed as a way to further examine the relation between ‘small’ collision systems and ‘large’ ones Brewer et al. (2021); Giacalone et al. (2025a, b). To this end, oxygen-oxygen (OO) collisions were performed at the Relativistic Heavy Ion Collider (RHIC) in May of 2021 at a center-of-mass collision energy per nucleon pair ($\sqrt{s_{\text{NN}}}$) of 200 GeV Huang (2023), and the Large Hadron Collider (LHC) performed OO and neon-neon (NeNe) collisions at $\sqrt{s_{\text{NN}}}=5.36$ TeV in July of 2025. The availability of these datasets and the first results coming from them CMS (2025a, b, c); Aad et al. (2025); Abualrob et al. (2025) have prompted intense interest in light ion collisions and, by extension, any experimental backgrounds that could affect measurements of observables attempting to gain insight into potential QGP-like effects that manifest in these collisions. Looking towards the future, light ion collisions have been proposed as a potential direction for the LHC ion physics program beyond the Run 4 era Waagaard et al. (2025); Citron et al. (2019), further motivating the need to study these backgrounds.
@@
Ion beams at a particle collider are subject to many effects that can cause beam losses which are not present for beams of more fundamental particles. As an example, for $^{208}\text{Pb}$ beams at the LHC, the conversion of a proton to a neutron via the process $^{208}\text{Pb}\to^{207}\text{Pb}+n$ causes the ion’s charge to mass ratio, which to very good approximation is equal to the atomic number $Z$ divided by the mass number $A$ when nuclear binding energies are neglected, to change. The daughter ion then has an incorrect magnetic rigidity for the the LHC beam optics and is unable to continue to circulate in the collider. This process accounts for a sizable fraction of Pb beam losses at the LHC Bruce et al. (2009).
@@
When encountering the strong electromagnetic fields caused by other ions, light ions are subject to electromagnetic dissociation processes Svetlichnyi et al. (2023); Pshenichnov et al. (1998) which can result in the ion breaking up. Examples of relevant processes for $^{16}\text{O}$ ions include $^{16}\text{O}\to^{14}\text{N}+\text{d}$, $^{16}\text{O}\to^{12}\text{C}+^{4}\text{He}$, and $^{16}\text{O}\to^{10}\text{B}+^{6}\text{Li}$. Unlike the scenario for heavy ions, the daughters of all of these processes have charge to mass ratios $Z/A=0.5$ that are identical to that of $^{16}\text{O}$ itself and are therefore able to continue to circulate indefinitely in the collider alongside the oxygen. This may result in beam contamination that grows over time as daughter ion species build up. For oxygen and neon, it is expected that $^{4}\text{He}$ will be the most likely contaminant due to its strong binding energy and the potential tendency for the nuclear structure of the original ion to contain $\alpha$ clusters Summerfield et al. (2021); Wang et al. (2024). Furthermore, neon may be more susceptible to dissociation, and therefore contamination effects, owing to its highly deformed nuclear structure Giacalone et al. (2025b). Electromagnetic dissociation is expected to occur most often at locations where ion beams are highly focused, i.e., at the collider’s interaction points where experiments sit, and should scale with the intensity of the beams at these locations. Thus, beam contamination effects may become a barrier preventing higher instantaneous luminosities in the future.
@@ II Method overview
The problem can be tackled by using an approach essentially equivalent to the ABCD method detailed in Ref Choi and Oh (2021); Kasieczka et al. (2021). In that method, two uncorrelated variables are utilized to define control regions to constrain backgrounds in order to search for signal in an adjacent region. Here, we utilize the expected time-dependence of the transmutation effect, alongside with another variable that is correlated with the “size” of the collision system as our two variables of interest. This second variable should be correlated with the number of interacting nucleons present in the collision, such that more ‘head-on’ collisions, or collisions of heavier ions, produce larger values. Examples of appropriate and common variables are the total number of tracks measured in the event $N_{\text{trk}}$ or the total energy observed in an experiment’s calorimeter system. For convenience of explanation, in the discussion that follows we choose to focus on using $N_{\text{trk}}$ as the second variable of choice. The proposed method utilizes the two-dimensional plane shown in Fig. 1, where the $N_{\text{trk}}$ distribution is plotted against the amount of time that has elapsed in a fill.
At the start of the fill, during a short a time interval that can be defined to be from times $t_{0}$ to $t_{1}$, the contribution from beam transmutation is expected to be small because there has not been sufficient time for contaminants to build up. We label this time as the reference control region in Figure 1. The reference control region can be used to establish the uncontaminated shape of the $N_{\text{trk}}$ distribution. Crucially, the shape of this distribution will not change versus fill time for non-contaminant collisions. However, one cannot simply subtract this uncontaminated distribution from distributions measured at times $t>t_{1}$, as the total rate of non-contaminated collisions is expected to be smaller for $t>t_{1}$ because of beam intensity decay.
@@
A scaling factor can be defined to account for the change overall scale in the non-contaminant distribution in a later time period $t_{1}<t<t_{2}$ as compared to the reference control region. It compares the uncontaminated tails of the $N_{\mathrm{trk}}$ distributions above $N_{\mathrm{trk}}^{\mathrm{cut}}$:
@@
Using this definition, one can extract the time-dependent $N_{\mathrm{trk}}$ distribution of the contaminant collisions ($N_{\mathrm{trk}}^{\text{cont.}}$) in the contaminated region of interest (which is defined as the space having $t>t_{1}$ and $N_{\mathrm{trk}}<N_{\mathrm{trk}}^{\mathrm{cut}}$, as seen in Fig. 1) by simply subtracting the scaled reference $N_{\mathrm{trk}}$ distribution:
@@ III Simulation studies
To test the validity of the proposed approach, a simple toy model study was performed using simulated events. In this model, beams of $^{16}\text{O}^{16}\text{O}$ collisions with only one contaminant species ($^{4}\text{He}$) are considered. Furthermore, contaminant-contaminant collisions ($^{4}\text{He}^{4}\text{He}$) are neglected, as they will be rare as compared to $^{4}\text{He}^{16}\text{O}$ collisions. Previously performed accelerator simulations for $^{16}\text{O}$ beams indicate that this approximation of only considering $^{4}\text{He}$ should capture the majority of the beam transmutation effects at the LHC Jowett et al. (2025).
To try to have the model more accurately reflect a realistic LHC fill, the time-dependent shape of the instantaneous $^{16}\text{O}^{16}\text{O}$ and $^{4}\text{He}^{16}\text{O}$ luminosity is extracted by performing a fit to the luminosity projections in Ref Jowett et al. (2025). Specifically, the time dependence is modeled with simple exponential functions: an exponential decay for OO collisions,
@@
$\frac{dN_{\mathrm{HeO}}}{dt}=a\,\left[1-\exp\!\left(-\frac{t}{\tau_{\mathrm{HeO}}}\right)\right],$ (4)
where $\tau_{\mathrm{HeO}}=8.1$ hours. Figure 2 shows the time evolution of the resulting event rates, $dN/dt$, up to 14 hours. A hadronic cross sections of 1.3b for OO collisions was used to convert the extracted instantaneous luminosity to an event rate. The event rate of contaminant HeO collisions, set by the $a$ parameter, is scaled to give an arbitrary 5% HeO contamination level by the end of a 14-hour fill. Note that the absolutes rates of OO and HeO collisions do not matter for the study; only their relative contributions matter for the purposes of illustrating the proposed technique.
Figure 2: The event rate, $dN/dt$, as a function of time up to 14 hours for OO and HeO collisions. The shapes of these distributions were extracted from Ref. Jowett et al. (2025).
Figure 3: The input $N_{\mathrm{trk}}$ distributions for OO and HeO events at t=0.0 hours (left), t=0.5 hours (middle) and t=14 hours (right), generated using HG-Pythia.
@@
To model the $N_{\mathrm{trk}}$ distributions for these collisions, we use the HG-Pythia model Loizides and Morsch (2017), a simple toy model that uses Glauber model for the initial state and Pythia for the final state. For a given MC Glauber event, PYTHIA (v6.28 with Perugia 2011 tune) is run to generate pp events with exactly ${{(N^{\mathrm{hard}}_{\mathrm{NN}}})}_{i}$ multiple parton interactions for each nucleon-nucleon collision $i$, where $i$ runs from 0 to $N_{\mathrm{coll}}$. The generated particles from all PYTHIA events are then combined and treated like a single AA event in the further analysis. HG-Pythia is not intended to necessarily correctly reproduce the overall scale OO $N_{\mathrm{trk}}$ distribution, but it does give a reasonable qualitative shape that can be used to illustrate the salient points of our proposed method. We generate both OO and HeO collisions at $\sqrt{s_{\mathrm{NN}}}$ = 5.36 TeV.
Figure 3 shows the input $N_{\mathrm{trk}}$ distributions used for three time slices, t=0 hours (left), t = 0.5 hours (middle) and t = 14 hours (right). These distributions are scaled to match the respective event rates $dN/dt$ from Figure 2. The $N_{\mathrm{trk}}$ distribution for OO is shown in open red markers and HeO is shown in open green markers. The sum of these distributions (which is what would actually be measured by an experiment) is shown by the blue line. The time-dependence of both of these distributions is evident after comparing the three panels: the OO collisions decay over time and the HeO collisions grow with time. This leads to a subtle change in shape of the blue line with time. Importantly, the input HeO distribution contributes significantly only in the low-$N_{\mathrm{trk}}$ region, approximately below $N_{\mathrm{trk}}$ = 900. Because of this, $N_{\mathrm{trk}}^{\mathrm{cut}}$ is chosen to be 900 for this implementation of our proposed method.
Figure 4 shows the results after applying our contamination extraction method. The blue lines are the same as in Fig. 3. From these total distributions, the OO and HeO contributions are extracted. The left panel shows the measurement in the reference control region, where the total is defined to be entirely resulting from OO collisions. The middle and right panels show how scaling the OO distribution using the high purity control region (illustrated by the red points), the difference between the total and the OO estimation gives a distribution for the HeO contribution (green points). The shapes of the output HeO $N_{\mathrm{trk}}$ distributions qualitatively match the shapes of the input distributions shown in 3. Note that small fluctuations in the range of $N_{\mathrm{trk}}\geq N_{\mathrm{trk}}^{\mathrm{cut}}$ are the result of statistical fluctuations when sampling the high $N_{\mathrm{trk}}$ tail of the total distribution.
@@
The validity of the method is assessed through a closure test, as shown in Figure 5. The closure is tested by comparing the extracted OO and HeO rates, calculated by integrating the distributions in Fig. 4, against the original inputs. Good agreement across most of the fill time demonstrates the ability to obtain reliable contamination estimates in this toy model. The deviations at low times can be attributed to statistical fluctuations from event sampling combined with the division by a very small input value since the contamination rate is very low at small times.

### v2

@@ Abstract
Collisions of relativistic light ions such as oxygen, neon, and magnesium, have been proposed as a way to examine the system-size dependence of dynamics typically associated with the quark-gluon plasma produced in collisions of heavier ions such as xenon, gold, or lead. Recent efforts at both the Relativistic Heavy Ion Collider (RHIC) and Large Hadron Collider (LHC) have produced large datasets of proton-oxygen, oxygen-oxygen, and neon-neon collisions, catalyzing intense interest in experimental backgrounds associated with light ion collisions. In particular, electromagnetic dissociation of light ions while they are circulating in a collider can result in beam contamination that is difficult to simulate precisely. Here we propose a data-driven method for evaluating the potential impact of beam contaminants on physics analyses. The method exploits the time-dependence and smaller size of contaminant ion species to define control regions that can be used to quantify potential contamination effects. A simple model is used to illustrate the method and to study its robustness. This method can inform studies of recent LHC and RHIC data and could also be useful for future light ion programs at the LHC and beyond.
@@ I Introduction
In the field of high energy nuclear physics, collisions of relativistic heavy ions, e.g., xenon, gold, or lead, are used to study the emergent properties of the strong force. In these collisions a strongly interacting hot medium, known as quark-gluon plasma (QGP), is created Busza:2018rrf; Niida:2021wut; Harris:2023tti. The QGP exhibits many interesting and nontrivial dynamics including collective fluid-like behavior Romatschke:2017ejr; Heinz:2013th and results in strong modifications to particle production as compared to proton-proton collision baselines CMS:2016uxf; ATLAS:2014ipv; ALICE:2013dpt.
@@
Perhaps surprisingly, critical examination of other collision systems having a much smaller size, such as proton-gold, deuteron-gold, helium-3-gold PHENIX:2017xrm, and proton-lead collisions ATLAS:2016yzd; CMS:2015yux have revealed similar behaviors as heavy ion collisions, prompting the question as to whether a QGP droplet is formed in these collisions. To try to gain more insight into this question, collisions of light ions such as oxygen, neon, and magnesium, have been proposed as a way to further examine the relation between ‘small’ collision systems and ‘large’ ones Brewer:2021kiv; Giacalone:2024ixe; Giacalone:2024luz. To this end, oxygen-oxygen (OO) collisions were performed at the Relativistic Heavy Ion Collider (RHIC) in May of 2021 at a center-of-mass collision energy per nucleon pair ($\sqrt{s_{{}_{\text{NN}}}}$) of 200 GeV Huang:2023viw, and the Large Hadron Collider (LHC) performed OO and neon-neon (NeNe) collisions at $\sqrt{s_{{}_{\text{NN}}}}=5.36$ TeV in July of 2025. The availability of these datasets and the first results coming from them CMS:2025bta; CMS:2025tga; ATLAS:2025nnt; ALICE:2025luc; CMS-PAS-HIN-25-010; CMS-PAS-HIN-25-014 have prompted intense interest in light ion collisions and, by extension, any experimental backgrounds that could affect measurements attempting to gain insight into potential QGP-like effects that manifest in these collisions. Looking towards the future, light ion collisions have been proposed as a potential direction for the LHC ion physics program beyond the Run 4 era Waagaard:2025dsi; Citron:2018lsq, further motivating the need to study these backgrounds.
@@
Ion beams at a particle collider are subject to many effects that can cause beam losses which are not present for beams of more fundamental particles. As an example, for $^{208}\text{Pb}$ beams at the LHC, the release of a neutron via the process $^{208}\text{Pb}\to{}^{207}\text{Pb}+n$ causes the ion’s charge to mass ratio, which to very good approximation is equal to the atomic number $Z$ divided by the mass number $A$ when nuclear binding energies are neglected, to change. The daughter ion then has an incorrect magnetic rigidity for the LHC beam optics and is unable to continue to circulate in the collider. This process accounts for a sizable fraction of Pb beam losses at the LHC Bruce:2009bg.
@@
When encountering the strong electromagnetic fields caused by other ions, light ions are subject to electromagnetic dissociation processes Svetlichnyi:2023nim; Pshenichnov:1997un which can result in the ion breaking up. Examples of relevant processes for $^{16}\text{O}$ ions include $^{16}\text{O}\to{}^{14}\text{N}+\text{d}$, $^{16}\text{O}\to{}^{12}\text{C}+^{4}\text{He}$, and $^{16}\text{O}\to{}^{10}\text{B}+{}^{6}\text{Li}$. Unlike the scenario for heavy ions, the daughters of all of these processes have charge to mass ratios $Z/A=0.5$ that are identical to that of $^{16}\text{O}$ itself and are therefore able to continue to circulate indefinitely in the collider alongside the oxygen. This may result in beam contamination that grows over time as daughter ion species build up. For oxygen and neon, it is expected that $^{4}\text{He}$ will be the most likely contaminant due to its strong binding energy and the potential tendency for the nuclear structure of the original ion to contain $\alpha$ clusters Summerfield:2021oex; YuanyuanWang:2024sgp. Furthermore, neon may be more susceptible to dissociation, and therefore contamination effects, owing to its highly deformed nuclear structure Giacalone:2024luz. Electromagnetic dissociation is expected to occur most often at locations where ion beams are highly focused, i.e., at the collider’s interaction points where experiments sit, and should scale with the intensity of the beams at these locations. Thus, beam contamination effects may become a barrier preventing higher instantaneous luminosities in the future.
@@ II Method overview
The problem can be tackled by using an approach essentially equivalent to the ABCD method detailed in Refs. Choi:2019mip; Kasieczka:2020pil. In that method, two uncorrelated variables are utilized to define control regions to constrain backgrounds in order to search for signal in an adjacent region. Here, we utilize the expected time-dependence of the transmutation effect, alongside with another variable that is correlated with the “size” of the collision system as our two variables of interest. This second variable should be correlated with the number of interacting nucleons present in the collision, such that more ‘head-on’ collisions, or collisions of heavier ions, produce larger values. Examples of appropriate and common variables are the total number of tracks measured in the event $N_{\text{trk}}$ or the total energy observed in an experiment’s calorimeter system. For convenience of explanation, in the discussion that follows we choose to focus on using $N_{\text{trk}}$ as the second variable of choice. The proposed method utilizes the two-dimensional plane shown in Figure 1, where the $N_{\text{trk}}$ distribution is plotted against the amount of time that has elapsed in a fill.
At the start of the fill, during a short a time interval that can be defined to be from times $t_{0}$ to $t_{1}$, the contribution from beam transmutation is expected to be small because there has not been sufficient time for contaminants to build up. We label this time as the reference control region in Fig. 1. The reference control region can be used to establish the uncontaminated shape of the $N_{\text{trk}}$ distribution. Crucially, the shape of this distribution will not change versus fill time for non-contaminant collisions. However, one cannot simply subtract this uncontaminated distribution from distributions measured at times $t>t_{1}$, as the total rate of non-contaminated collisions is expected to be smaller for $t>t_{1}$ because of beam intensity decay.
@@
A scaling factor, $f$, can be defined to account for the change in the overall scale of the non-contaminant distribution in a later time period $t_{1}<t<t_{2}$ as compared to the reference control region. It compares the uncontaminated tails of the $N_{\mathrm{trk}}$ distributions above $N_{\mathrm{trk}}^{\mathrm{cut}}$:
@@
Using this definition, one can extract the time-dependent $N_{\mathrm{trk}}$ distribution of the contaminant collisions, $N_{\mathrm{trk}}^{\text{cont.}}$, in the contaminated region of interest (which is defined as the space having $t>t_{1}$ and $N_{\mathrm{trk}}<N_{\mathrm{trk}}^{\mathrm{cut}}$, as seen in Fig. 1) by simply subtracting the scaled reference $N_{\mathrm{trk}}$ distribution:
@@ III Simulation studies
To test the validity of the proposed approach, a simple toy model study was performed using simulated events. In this model, beams of $^{16}\text{O}^{16}\text{O}$ collisions with only one contaminant species ($^{4}\text{He}$) are considered. Furthermore, contaminant-contaminant collisions ($^{4}\text{He}^{4}\text{He}$) are neglected, as they will be rare as compared to $^{4}\text{He}^{16}\text{O}$ collisions. Previously performed accelerator simulations for $^{16}\text{O}$ beams indicate that this approximation of only considering $^{4}\text{He}$ should capture the majority of the beam transmutation effects at the LHC transmutation.
To try to have the model more accurately reflect a realistic LHC fill, the time-dependent shape of the instantaneous $^{16}\text{O}^{16}\text{O}$ and $^{4}\text{He}^{16}\text{O}$ luminosity is extracted by performing a fit to the luminosity projections in Ref transmutation. Specifically, the time dependence is modeled with simple exponential functions: an exponential decay for OO collisions,
@@
$\frac{dN_{\mathrm{HeO}}}{dt}=b\,\left[1-\exp\!\left(-\frac{t}{\tau_{\mathrm{HeO}}}\right)\right],$ (4)
where $\tau_{\mathrm{HeO}}=8.1$ hours. Figure 2 shows the time evolution of the resulting event rates, $\mathrm{d}N/\mathrm{d}t$, up to 14 hours. A hadronic cross section of 1.3b for OO collisions was used to convert the extracted instantaneous luminosity to an event rate, which defines the overall scale factor $a$. The event rate of contaminant HeO collisions, set by the $b$ parameter, is scaled to give an arbitrary 5% HeO contamination level by the end of a 14-hour fill. Note that the absolutes rates of OO and HeO collisions do not matter for the study; only their relative contributions matter for the purposes of illustrating the proposed technique.
Figure 2: The event rate, $\mathrm{d}N/\mathrm{d}t$, as a function of time up to 14 hours for OO and HeO collisions. The shapes of these distributions were extracted from Ref. transmutation.
Figure 3: The input $N_{\mathrm{trk}}$ distributions for OO and HeO events at $t=0$ hours (left), $t=0.5$ hours (middle) and $t=14$ hours (right), generated using HG-Pythia.
@@
To model the $N_{\mathrm{trk}}$ distributions for these collisions, we use the HG-Pythia model Loizides:2017sqq, a simple toy model that uses Glauber model for the initial state and Pythia for the final state. For a given MC Glauber event, PYTHIA (v6.28 with Perugia 2011 tune) is run to generate pp events with exactly ${{(N^{\mathrm{hard}}_{\mathrm{NN}}})}_{i}$ multiple parton interactions for each nucleon-nucleon collision $i$, where $i$ runs from 0 to the total number of nucleon-nucleon collisions, $N_{\mathrm{coll}}$. The generated particles from all PYTHIA events are then combined and treated like a single ion collision in the further analysis. HG-Pythia is not intended to necessarily correctly reproduce the overall scale of the OO $N_{\mathrm{trk}}$ distribution, but it does give a reasonable qualitative shape that can be used to illustrate the salient points of our proposed method. We generate both OO and HeO collisions at $\sqrt{s_{\mathrm{{}_{\text{NN}}}}}$ = 5.36 TeV.
Figure 3 shows the input $N_{\mathrm{trk}}$ distributions used for three time slices, t=0 hours (left), t = 0.5 hours (middle) and t = 14 hours (right). These distributions are scaled to match the respective event rates $\mathrm{d}N/\mathrm{d}t$ from Fig. 2. The $N_{\mathrm{trk}}$ distribution for OO is shown in red markers and HeO is shown in green markers. The sum of these distributions (which is what would actually be measured by an experiment) is shown by the black line. The time-dependence of both of these distributions is evident after comparing the three panels: the OO collisions decay over time and the HeO collisions grow with time. This leads to a subtle change in shape of the black line with time. Importantly, the input HeO distribution contributes significantly only in the low-$N_{\mathrm{trk}}$ region, approximately below $N_{\mathrm{trk}}=900$. Because of this, $N_{\mathrm{trk}}^{\mathrm{cut}}$ is chosen to be $900$ for this implementation of our proposed method.
Figure 4 shows the results after applying our contamination extraction method. The black lines are the same as in Fig. 3. From these total distributions, the OO and HeO contributions are extracted. The left panel shows the measurement in the reference control region, where the total is defined to be entirely resulting from OO collisions. The middle and right panels show how scaling the OO distribution using the high purity control region (illustrated by the red markers), the difference between the total and the OO estimation gives a distribution for the HeO contribution (green markers). The shapes of the output HeO $N_{\mathrm{trk}}$ distributions qualitatively match the shapes of the input distributions shown in Fig. 3. Note that small fluctuations in the range of $N_{\mathrm{trk}}\sim N_{\mathrm{trk}}^{\mathrm{cut}}$ are the result of statistical fluctuations when sampling the high $N_{\mathrm{trk}}$ tails of the distributions.
@@
The validity of the method is assessed through a closure test, as shown in Figure 5. The closure is tested by comparing the extracted OO and HeO rates, calculated by integrating the distributions in Fig. 4, against the original inputs. The sub-percent-level agreement observed across the fill time demonstrates the ability of the method to provide reliable contamination estimates in this toy model.